Imagine you are standing 4 meters away from a painting hung on a wall. If the top of the painting is 3 meters from the ground, which expressions represent the angle, θ, that your feet make with the top of the painting?
Answer by Boreal(15235)

tangent of theta=3/4
sin theta=3/5, opposite over hypotenuse, the 5 comes from the fact it is 5 meters from the feet to the top of the painting. This is a 3-4-5 right triangle
cos theta=4/5, adjacent over hypotenuse